# Exploring Families’ Use and Mediation of Generative AI: A Multi-User Perspective

### v2

@@ 6. Discussion
We investigated how the perceptions of GenAI and contextual characteristics of families contributed to their GenAI mediation strategies. In our findings, we proposed an adapted version of parental media mediation literature in the context of Family-GenAI interactions. We then applied this model to describe six observed Family-GenAI mediation and use profiles. Overall, our findings suggest that although most parents used at least one type of mediation strategy to manage their children’s interactions with ChatGPT and similar GenAI, they did not always utilize the same mediation strategies they had for other media their children accessed due to varying perspectives on or knowledge about GenAI. GenAI was often viewed as different from existing media (e.g., social media, television, or video platforms), therefore lending itself to different mediation strategies. Our results carry design implications for GenAI platform providers and research implications for broadening existing parental media mediation frameworks.
It should be noted that these findings, as well as several of our profiles, align with Rubin et al.’s recent publication (Rubin et al., 2025), which also presented profiles to describe how parents’ perceptions of GenAI affected how they use GenAI with their families. Though our research took a slightly different approach in its inclusion of children’s perspectives within each family, we consider our results to be complementary to Rubin et al.’s.
@@ 6.1.1. Existing online safety rules are insufficient for GenAI
Although most of the participating families reported adapting some mediation strategies they used for other online media (e.g., TV media, news outlets) to GenAI, GenAI may have unique nuances that need careful attention. Specifically, GenAI is interactive and mimics human capabilities, but is incapable of human-like thinking. While these attributes create opportunities for positive user outcomes, GenAI lacks a governance framework and generates information that might be incomplete, incorrect, and difficult to validate. These nuances require families to learn new ways to mediate this fast-evolving media source. We anticipate that GenAI-tailored mediation strategies will evolve with time and experience. For example, parents or educators might experiment with new mediation strategies for GenAI as they learn more about the technology. However, they may need support in learning how to develop or tailor mediation strategies to GenAI’s unique affordances.
@@
Parents in our study expressed concerns about the safety of other interactive technologies, such as social media and voice assistants. Half of the parents worried about their children being addicted to social media or being spied on via phones or voice assistants, and eight families expressed greater trust in GenAI than other technologies and media. This concern misalignment may be because GenAI is a relatively recent technology that families have limited experience with. As a result, families may still be shaping their mental models of this technology. Since we recruited families who used GenAI on shared accounts, they may have been more trusting of the platform than families not on GenAI platforms. Concerns about ChatGPT or GenAI were more similar to their concerns about search engines (i.e., Google) than social media, suggesting ChatGPT is more often perceived as an information resource rather than a social or entertainment platform. For example, the parent in F4 shared an analogy of how their generation learned to shift from encyclopedias to search engines, similar to how children nowadays substitute search engines with ChatGPT. However, ChatGPT is a language model trained on existing data, and users were not clearly informed of this. However, the assumption that ChatGPT produces search-engine-quality information could create vulnerabilities if families become overly reliant on GenAI. Parents’ trust in ChatGPT might also be based on their awareness of the built-in guardrails. Even families who recognized that such mechanisms could be bypassed still felt ChatGPT was safer than social media. However, these guardrails may not be sufficiently comprehensive  (Liu et al., 2023) and are not tailored to children’s and families’ use.
@@ 6.1.3. Participating parents did not know how to tailor existing mediation strategies to GenAI
In reflecting on GenAI, participants were most concerned about the technology (1) violating their children’s privacy, (2) negatively influencing behavior and learning (e.g., over-reliance, cheating), (3) generating inappropriate content or misinformation (consistent with Yu et al. (2024)). The strategies parents used to mitigate these risks were often related to managing their children’s behaviors and practices, including not revealing private information, having family conversations around safe use of the Internet, practicing how to validate information. No families implemented any ChatGPT-specific technical safety mediation strategies similar to what they might do on other websites (e.g., router-level block for certain websites). In this context, technical safety (Nikken and Jansz, 2014) refers to mediation implemented on applications on the devices, such as a blocklist filter to mediate children’s website access. We consider two possible explanations for this: (1) not knowing what technical mediation to implement and (2) how to implement them. First, GenAI has unique affordances that distinguish it from other media, which introduces new risks and opportunities. It is possible that parents may have difficulty identifying the appropriate technical tools to manage ChatGPT use. Furthermore, parents who have existing technical tools may be unsure of how to tailor them to ChatGPT. Although families may develop or adapt mediation strategies with time and experience, limited resources are available for parents hoping to implement GenAI-specific technical safety measures. These observations highlight a gap in learning resources tailored to the Family-GenAI context that should be addressed in future work in this field.
@@ 6.1.4. Co-use can potentially support family bonding
A promising and positive observation from this study was the number of families that co-used ChatGPT together. Many families shared reasons for co-using AI to spend time together, including solving homework problems (F5), editing their writing (F11), disputing debates (F2, F5), and generating and sharing stories with one another (F7, F8, F12). We also observed at least one family (F6) who assigned ChatGPT a facilitating ‘role’ during their game nights. Notably, being able to use ChatGPT allowed the family, which only consisted of a parent and a child, to try games that they would otherwise need a third person to play. Overall, there are opportunities for future work to explore emergent and novel joint-media engagement specific to GenAI.
@@ 6.2. Design Implications – Platform Providers and Developers
Here, we translate our aforementioned reflections into practical design implications for platform providers and developers.
@@
  • Empower Families: Platforms should offer parents options to implement their desired mediation strategies. For instance, the interface could report children’s chatbot queries in real time or notify parents if personal information is shared with the chatbot. Platforms could offer profile options meant for family use that provide ways for families to tailor based on their concerns, use cases, desired controls, etc. To promote co-use, platforms may consider providing prompts that encourage family interactions (e.g., game ideas) and conversations (e.g., about online safety and misinformation).
Figure 3. Illustration of how family profiles (in Section 5) are situated with respect to these three dimensions: Control over GenAI usage, Trust over GenAI, and Co-Use toward GenAI.
[…]
A graph showing six family profiles’ three dimensions: Control over GenAI usage, Trust over GenAI, and Co-Use toward GenAI. The X-axis represents Trust, and the Y-axis represents Control. Both axes have three levels: low, medium, and high. The data dot’s size represents their co-use frequency, with small being less co-use. Skeptical family is located at the low trust and high control position, with a medium dot representing medium co-use. Hands-On family is located at medium trust and high control position, with a medium dot representing medium co-use. Use together family is located at medium trust and high control position, with a large dot representing more co-use. Cautious family is located at high trust and high control position, with a small dot representing less co-use. Trusting family is located at high trust and medium control position, with a medium dot representing medium co-use. Independent family is located at the medium trust and low control position, with a small dot representing less co-use.
@@ 6.3. Limitations and Future Work
First, our small sample size (N=12), while typical of qualitative research, limits the generalizability of our findings. We also specifically targeted families who shared a ChatGPT account and recruited participants via social media advertisements. While this meant we could recruit participants beyond our local area, it also constrained our sample size due to challenges in screening spam entries and identifying eligible families. Furthermore, our participant pool likely reflects a bias toward technology-savvy families who already use social media and GenAI, resulting in higher-than-average education levels and potentially greater trust in the technology. Future work could focus more on families who are less familiar with technology in general or are more hesitant about using ChatGPT, as they might employ different mediation strategies. Including families that do not use ChatGPT would further diversify the sample and provide broader insights.
Second, our study focused primarily on ChatGPT. Although we welcomed insights on other LLM-based GenAI tools (e.g., Deepseek, Gemini), our recruitment efforts were tailored to ChatGPT, as it was the most commonly used platform during the time of recruitment.
Third, interviewing children and parents together may have influenced participants’ responses. While parental presence can help children provide more detailed answers, it might also limit or shape what children are willing to share  (Gardner and Randall, 2012). Interviewing family members both together and individually could provide deeper insights into inter-family concerns, such as the shared use of a conversational agent capable of storing prior interactions.
@@ 7. Conclusion
This study explored families’ use and co-use experiences on text-based GenAI platforms (i.e., ChatGPT). We identified Family-AI mediation and use scenarios and proposed a model to describe how families determine the types of mediation strategies they use, and inversely, how these strategies may affect their perceptions of GenAI. Additionally, we identified six family profiles (i.e., skeptical, cautious, hands-on, use together, trusting, and independent) with variations in parents’ level of control over their children’s GenAI use and families’ trust toward and co-usage of GenAI. Our work highlights the unique risks and opportunities families with children face in their interactions with GenAI, and offers considerations for GenAI designers and developers interested in creating platforms that are both safe and valuable for children and their parents.

### v3

@@ 5. Discussion
We investigated how the perceptions of GenAI and contextual characteristics of families contributed to their GenAI mediation strategies. In our findings, we introduce the emerging themes from thematic analysis. We then proposed an adapted version of parental media mediation literature in the context of Family-GenAI interactions. Finally, we describe family GenAI interactions through the lens of our developed model. Overall, our findings suggest that although most parents used at least one type of mediation strategy to manage their children’s interactions with ChatGPT and similar GenAI, they did not always utilize the same mediation strategies they had for other media their children accessed due to varying perspectives on or knowledge about GenAI. GenAI was often viewed as different from existing media (e.g., social media, television, or video platforms), therefore lending itself to different mediation strategies. Our results carry design implications for GenAI platform providers and research implications for broadening existing parental media mediation frameworks.
@@ 5.1.1. Existing online safety rules are insufficient for GenAI
Although most of the participating families reported adapting some mediation strategies they used for other online media (e.g., TV media, news outlets) to GenAI, GenAI may have unique nuances that need careful attention. Specifically, GenAI is interactive and mimics human capabilities, but is incapable of human-like thinking. While these attributes create opportunities for positive user outcomes, GenAI lacks a governance framework and generates information that might be incomplete, incorrect, and difficult to validate. (taeihagh2025governance; huang2025survey; bashardoust2024comparing) These nuances require families to learn new ways to mediate this fast-evolving media source. We anticipate that GenAI-tailored mediation strategies will evolve with time and experience. For example, parents or educators might experiment with new mediation strategies for GenAI as they learn more about the technology. However, they may need support in learning how to develop or tailor mediation strategies to GenAI’s unique affordances.
@@
Parents in our study expressed concerns about the safety of other interactive technologies, such as social media and voice assistants. Half of the parents worried about their children being addicted to social media or being spied on via phones or voice assistants, and eight families expressed greater trust in GenAI than other technologies and media. This concern of misalignment may be because GenAI is a relatively recent technology that families have limited experience with. As a result, families may still be shaping their mental models of this technology. Concerns about ChatGPT or GenAI were more similar to their concerns about search engines (i.e., Google) than social media, suggesting ChatGPT is more often perceived as an information resource rather than a social or entertainment platform. For example, the parent in F4 shared an analogy of how their generation learned to shift from encyclopedias to search engines, similar to how children presently substitute search engines with ChatGPT. However, ChatGPT is a language model trained on existing data, and users were not clearly informed of this. However, the assumption that ChatGPT produces search-engine-quality information could create vulnerabilities if families become overly reliant on GenAI. Parents’ trust in ChatGPT might also be based on their awareness of the built-in guardrails. Even families who recognized that such mechanisms could be bypassed still felt ChatGPT was safer than social media. However, these guardrails may not be sufficiently comprehensive  (liu2023jailbreaking) and are not tailored to children’s and families’ use.
@@ 5.1.3. Participating parents did not know how to tailor existing mediation strategies to GenAI
In reflecting on GenAI, participants were most concerned about the technology (1) violating their children’s privacy, (2) negatively influencing behavior and learning (e.g., over-reliance and cheating), (3) generating inappropriate content or misinformation (consistent with yu2024exploring). The strategies parents used to mitigate these risks were often related to managing their children’s behaviors and practices, including not revealing private information, having family conversations around safe use of the Internet, and practicing how to validate information. No families implemented any ChatGPT-specific technical safety mediation strategies similar to what they might do on other websites (e.g., router-level block for certain websites). In this context, technical safety (nikken2014developing) refers to mediation implemented on applications on the devices, such as a blocklist filter to mediate children’s website access. We offer two possible explanations for this: (1) not knowing what technical mediation to implement and (2) how to implement them. First, GenAI has unique affordances that distinguish it from other media, which introduce new risks and opportunities. It is possible that parents may have difficulty identifying the appropriate technical tools to manage ChatGPT use. Furthermore, parents who have existing technical tools may be unsure of how to tailor them to ChatGPT. Although families may develop or adapt mediation strategies with time and experience, limited resources are available for parents hoping to implement GenAI-specific technical safety measures. These observations highlight a gap in learning resources tailored to the Family-GenAI context that should be addressed in future work in this field.
@@ 5.1.4. Co-use can potentially support family bonding
A promising and positive observation from this study was the number of families that co-used ChatGPT together. Many families shared reasons for co-using AI to spend time together, including solving homework problems (F5), editing their writing (F11), disputing debates (F2, F5), and generating and sharing stories with one another (F7, F8, F12). We also observed at least one family (F6) that assigned ChatGPT a facilitating ‘role’ during their game nights. Notably, being able to use ChatGPT allowed the family, which only consisted of a parent and a child, to try games that would otherwise need a third person to play. Overall, there are opportunities for future work to explore emergent and novel joint-media engagement specific to GenAI.
@@ 5.2. Design Implications – Platform Providers and Developers
Here, we translate our reflections into practical design implications for platform providers and developers.
@@
  • Empower Families: Platforms should offer parents options to implement their desired mediation strategies. For instance, the interface could report children’s chatbot queries in real time or notify parents if personal information is shared with the chatbot. Platforms could offer profile options meant for family use that provide ways for families to tailor based on their concerns, use cases, desired controls, etc.
@@ 5.3. Limitations and Future Work
First, the small sample size (N=12), while typical of qualitative research, limits the generalizability of our findings. The source dataset also specifically targeted families who shared a ChatGPT account and recruited participants via social media advertisements. Furthermore, the participant pool likely reflects a bias toward technologically-savvy families who already use social media and GenAI, resulting in higher-than-average education and income levels and potentially greater trust in the technology. Future work could focus more on families who are less familiar with technology in general or are more hesitant about using ChatGPT, as they might employ different mediation strategies. Including families that do not use ChatGPT would further diversify the sample and provide broader insights.
Second, the interviews have children and parents together all the time, which may have influenced participants’ responses. While parental presence can help children provide more detailed answers, it might also limit or shape what children are willing to share  (gardner2012effects). Interviewing family members both together and individually could provide deeper insights into inter-family concerns, such as the shared use of a conversational agent capable of storing prior interactions.
@@ 6. Conclusion
This study explored families’ use and co-use experiences on text-based GenAI platforms (i.e., ChatGPT). We identified Family-AI mediation and use scenarios and proposed a model to describe how families determine the types of mediation strategies they use, and inversely, how these strategies may affect their perceptions of GenAI. Additionally, we provide an in-depth presentation of the participating families’ GenAI interaction through our proposed model. Our work highlights the unique risks and opportunities families with children face in their interactions with GenAI, and offers considerations for GenAI designers and developers interested in creating platforms that are both safe and valuable for children and their parents.